OpenStudy (anonymous):

What is the solution to the system below? -3x-y=-3 9x+3y=9 (3, 3) (3, -3) no solution infinite solutions on the line

OpenStudy (anonymous):

explain to me how to do this like how to start this off

OpenStudy (anonymous):

Ok, so your possibilities are: Infinitely many solutions. No solutions. One solution. I put this in this order because you'll want to check each one on it's own.

OpenStudy (anonymous):

So first, to check to see if it is infinite solutions: If you have 2 equations (eq1 and eq2) If eq1 = k * eq2 (or vice versa) where k is any number, then you will have infinite solutions.

OpenStudy (anonymous):

Can you see a number which ,if you multiplied one of the equations by it, would give you the other equation?

OpenStudy (anonymous):

ok so i have to slove the equations first right??

OpenStudy (anonymous):

You can if you like.

OpenStudy (anonymous):

It's not really needed for 2 of the 3 cases.

OpenStudy (anonymous):

for example -3x-y=-3 -y=-3x-3

OpenStudy (anonymous):

That's not quite solved yet. Need to get rid of the - in front of the y.

myininaya (myininaya):

did you mean to add 3x on both sides?

OpenStudy (anonymous):

Assuming you're solving for y.

OpenStudy (anonymous):

Also yes, it should be a +3x

OpenStudy (anonymous):

yea m sloving for y

OpenStudy (anonymous):

Ok, so you add 3x to both sides you'll have: -y = 3x - 3

OpenStudy (anonymous):

Then multiply both sides by -1

OpenStudy (anonymous):

y=-3x-3

OpenStudy (anonymous):

y = -3x + 3

OpenStudy (anonymous):

You have to multiply -1 times both the 3x and the -3

OpenStudy (anonymous):

y i dnt get that y do i need to * -1??

OpenStudy (anonymous):

to get rid of the - in front of the y.

OpenStudy (anonymous):

ok i already did that o the y is now postive so its postive y=-3x-3

OpenStudy (anonymous):

\(-3x-y=-3\) +3x +3x \(\implies -y = 3x - 3\) *-1 *-1 \(\implies y = -3x + 3\)

OpenStudy (anonymous):

No, you have to multiply the whole right side by -1. \(-1(3x - 3) = -3x + 3\)

OpenStudy (anonymous):

ok i already did that o the y is now postive so its postive y=-3x-3

OpenStudy (anonymous):

yes that is wat i did soooo... wat do i do next do i slove the nxt equation?? qwhich would be y=-3x-3

OpenStudy (anonymous):

NO. It's y = -3x + 3

OpenStudy (anonymous):

ok i already did that o the y is now postive so its postive y=-3x-3

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

now do the same thing for the other equation.